Find the slope and y-intercept of the line 8x-3y=21
Answer by mananth(12270)

Find the slope and y-intercept of the line
8x-3y=21
3y=8x-21
/3
y= (8/3)x-7
compare with the equation y=mx+b
slope m = 8/3
y intercept = b = -7